OpenStudy (anonymous):

Solve using the square root method. Round to the nearest tenth, if necessary. x2 + 17 = 0 {10.7} {-4.1, 4.1} {-17, 17} No real solution

OpenStudy (anonymous):

sorry my being slow im in 9th trying my best to get to 10th

OpenStudy (phi):

you have \[ x^2 + 17 = 0 \] the first step is add -17 to both sides \[ x^2 + 17-17 = 0-17 \] and simplify

OpenStudy (phi):

Can you simplify \[ x^2 + 17-17 = 0-17 \] ? you should get \[ x^2 = -17\] now this is important: there is no (real) number times itself that will be negative

OpenStudy (phi):

normally you would take the square root of both sides \[ \sqrt{x^2} = \sqrt{-17} \\ x = \sqrt{-17} \] but x will not be a real number.

OpenStudy (phi):

if you type sq rt of -17 into your calculator, it will complain

OpenStudy (phi):

I assume you figured out the answer is "no real solution"